OpenStudy (ray_sihota):

having an issue finding the 3 point for 3x+5y=6 I get the frist 2 points 0, 6/5 and -2, 0 but having issue getting the thrid point

OpenStudy (anonymous):

put x=1 3x1+5y=6 y=3/5 (1, 3/5)

OpenStudy (anonymous):

put x=2 3x2+5y=6 y=0 (2, 0)

OpenStudy (ray_sihota):

then is there an easy way to find a 3rd pair

OpenStudy (anonymous):

nope it is easiest

OpenStudy (ray_sihota):

do not get it I am not getting a strait line

OpenStudy (ray_sihota):

I got it , thanks for your help

OpenStudy (anonymous):

:) happy to help